# No-Regret Strategy Solving in Imperfect-Information Games via Pre-Trained Embedding

### v1

@@ C.2 The Construction of the Hand Strength Tensor $\mathbf{y}_{q}$
In Numeral211 Hold’em, for example, during the terminal round, each player holds one hole card, and there is one community card on the table. For a specific hand $x$ (composed of the player’s hole card and the community card), the comparison set includes all other possible hands formed by pairing each of the remaining 38 hole cards with the same community card (i.e., $x^{\prime}$). The vector $\mathbf{w}$ is thus calculated based on how $x$ performs against these 38 potential opponent hands.
@@
C.3 Structure of the Hand Feature Encoder
@@
The hand feature encoder (HFE) is a simple structure. Since we model a hand as an image-like tensor, the HandEbdNet — essentially a regression task mapping $\mathbf{x}_{q}$ to $\mathbf{y}_{q}$ — adopts a CNN architecture to identify hand patterns (see Figure 7). Logically, HFE comprises a convolutional layer followed by a fully connected layer.
@@
C.4 Traning and Inference
We train a separate network for each round, considering only Round 2 and Round 3. As shown in Figure 7, we need to specify the “round” parameter. The training data consists of all possible hand combinations in the current round: for Round 2, the number of combinations is $\binom{40}{2}\times\binom{38}{1}=780\times 38=29,\!640$; for Round 3, it is $\binom{40}{2}\times\binom{38}{1}\times\binom{37}{1}=1,\!096,\!680$. Such data volumes are easily traversable and trainable on a home PC. Since the same data is used during inference, we do not need to consider issues of network robustness or overfitting — this is essentially a simple regression task, which is why a basic network structure suffices to solve this task.
Appendix D Experiment

### v2

@@ D.2 The Construction of the Hand Strength Tensor $\mathbf{y}_{q}$
In Numeral211 Hold’em, for example, during the terminal round, each player holds one private card, and there is one community card on the table. For a specific hand $x$ (composed of the player’s private card and the community card), the comparison set includes all other possible hands formed by pairing each of the remaining 38 private cards with the same community card (i.e., $x^{\prime}$). The vector $\mathbf{w}$ is thus calculated based on how $x$ performs against these 38 potential opponent hands.
@@
D.3 Structure of the Hand Feature Encoder
The hand feature encoder (HFE) is a simple structure. Since we model a hand as an image-like tensor, the HandEbdNet–essentially a regression task mapping $\mathbf{x}_{q}$ to $\mathbf{y}_{q}$–adopts a CNN architecture to identify hand patterns (see Figure 7). Logically, HFE comprises a convolutional layer followed by a fully connected layer.
@@ D.3 Structure of the Hand Feature Encoder
D.4 Training and Inference
We train a separate network for each round, considering only Round 2 and Round 3. As shown in Figure 7, we need to specify the “round” parameter. The training data consists of all possible hand combinations in the current round: for Round 2, the number of combinations is $\binom{40}{2}\times\binom{38}{1}=780\times 38=29,\!640$; for Round 3, it is $\binom{40}{2}\times\binom{38}{1}\times\binom{37}{1}=1,\!096,\!680$. Such data volumes are easily traversable and trainable on a home PC. Since the same data is used during inference, we do not need to consider issues of network robustness or overfitting–this is essentially a simple regression task, which is why a basic network structure suffices to solve this task.
Appendix E Experiment